# Enhancing the ODMR Signal of Organic Molecular Qubits

### v2

@@ Conclusion
By covalently tethering two benzylic radicals at their para positions, the resulting diradical of $D_{2}$ point group symmetry has a triplet GS of irrep $B_{1}$, lowest singlet CTs of irreps $A$ and $B_{1}$, and lowest triplet LEs of irreps $B_{2}$ and $B_{3}$. Thus, by group-theoretical considerations, spin polarisation can be achieved via a $\Delta M_{S}=\pm 1$ ISC from LEs to CTs, followed by a $\Delta M_{S}=0$ ISC from CTs to GSs, just like in an NV centre. Moreover, the associated SOC matrix elements are appreciable due to next-nearest-neighbour effects for the former and nearest-neighbour effects for the latter. We expect this success to be generalisable to any diradical based upon two para-connected benzylic radicals. This is illustrated in the PT2TM-p-PT2TM diradical: Firstly, the benzylic radicals are meta-chlorinated to lock the inter-radical torsion by steric repulsion, ensuring that the GS comprises both triplets and singlets at near-degeneracy. Next, the stabilising phenyl rings are specially chosen to not break the $C_{2}$ rotational symmetry down the benzylic ipso-para direction. Finally, every benzylic radical has one more chloro substituent than each of its phenyl branches, thereby localising the lowest-lying excitations on the tetra-chlorinated benzylic fragment. Indeed, these expectations are reflected in the ab initio results by producing the desired electronic structure for ODMR.
To our best knowledge, our new design has checked most boxes for a robust optical-spin interface [17]. Perhaps a minor problem not addressed by this work is the poor emissive properties of our PT2TM-p-PT2TM prototype due to its alternacy symmetry [93, 94, 51]. For that, we expect techniques of alternacy symmetry breaking [93, 94, 45] and excited-state symmetry breaking [59, 60] employed by the organic light-emitting diode (OLED) community to be useful. In fact, the latter was recently demonstrated by Chowdhury et al. [52], who synthesised trityl-based diradicals of near-unity luminescence quantum yields by mesitylating at the para positions. When applied to PT2TM-p-PT2TM, a possible diradical would look like Fig. Introductionc, which is a suitable synthetic starting point. Notably, the ODMR contrast for this novel spin polarisation pathway is expected to be around $30\%$, an unequivocal improvement from earlier works [51, 52, 53]. Regarding qubit operations, our design also overcomes problems of Zeeman-induced spin decoherence and observations of multiple Rabi frequencies during EPR measurements of weakly-coupled diradicals. These are all crucial steps towards the realisation of optically-addressable molecular spin qubits.

### v3

@@ Conclusion
By covalently tethering two benzylic radicals at their para positions, the resulting diradical of $D_{2}$ point group symmetry has a triplet GS of irrep $B_{1}$, lowest singlet CTs of irreps $A$ and $B_{1}$, and lowest triplet LEs of irreps $B_{2}$ and $B_{3}$. Thus, by group-theoretical considerations, spin polarisation can be achieved via a $\Delta M_{S}=\pm 1$ ISC from LEs to CTs, followed by a $\Delta M_{S}=0$ ISC from CTs to GSs, just like in an NV centre. Moreover, the associated SOC matrix elements are appreciable due to next-nearest-neighbour effects for the former and nearest-neighbour effects for the latter. We expect this success to be generalisable to any diradical based upon two para-connected benzylic radicals. This is illustrated in the PT2TM-p-PT2TM diradical: Firstly, the benzylic radicals are meta-chlorinated to lock the inter-radical torsion by steric repulsion, ensuring that the GS comprises both triplets and singlets at near-degeneracy. Next, the stabilising phenyl rings are specially chosen to not break the $C_{2}$ rotational symmetry down the benzylic ipso-para direction, which is the essential symmetry element of the aforementioned benzylic para-dimer. Finally, every benzylic radical has one more chloro substituent than each of its phenyl branches, thereby localising the lowest-lying excitations on the tetra-chlorinated benzylic fragment. Indeed, these expectations are reflected in the ab initio results by producing the desired electronic structure for ODMR.
To our best knowledge, our new design has checked most boxes for a robust optical-spin interface [17]. Perhaps a minor problem not addressed by this work is the poor emissive properties of our PT2TM-p-PT2TM prototype due to its alternacy symmetry [98, 99, 52]. For that, we expect techniques of alternacy symmetry breaking [98, 99, 45] and excited-state symmetry breaking [60, 61] employed by the organic light-emitting diode (OLED) community to be useful. In fact, the latter was recently demonstrated by Chowdhury et al. [53], who synthesised trityl-based diradicals of near-unity luminescence quantum yields by mesitylating at the para positions. When applied to PT2TM-p-PT2TM, a possible diradical would look like Fig. Introductionc, which is a suitable synthetic starting point. Notably, the ODMR contrast for this novel spin polarisation pathway is expected to be around $30\%$ with more than $25\%$ of the population being in the triplet manifold at steady state, an unequivocal improvement from earlier theoretical [52] and experimental [53, 54] works. Regarding qubit operations, our design also overcomes problems of Zeeman-induced spin decoherence and observations of multiple Rabi frequencies during EPR measurements of weakly-coupled diradicals. These are all crucial steps towards the realisation of optically-addressable molecular spin qubits.